OpenStudy (anonymous):

Hi I have an alfa particle with a KE o 850 eV, it has a mass of 6,68x10^-27 kg and a charge of 3,2x10^-19 C . How can I calculate the wavelength of it, from De Broglie equation. The particle itself is placed in uniform electric field.

OpenStudy (jamesj):

From the deBroglie relations, wavelength, lambda is given by \[ \lambda = h / p \] where h is the Planck constant and p is the momentum. You can calculate momentum (assuming the particle isn't moving at relativistic speed) by calculating from its KE, it's velocity v and hence p = mv.

OpenStudy (anonymous):

Ok Ill try this wait a sec

OpenStudy (anonymous):

I believe it is moving at a relativistic speed, because if I calculate the velocity as a square root of twice the kintic energy divided by the mass, it comes out to be in a factor of 5,04 x 10^14. Now that is more than twice the speed of light.

OpenStudy (jamesj):

In which case, it's a bit of a pain. Use this tool here, but make sure of course (!) you understand the formulae: http://hyperphysics.phy-astr.gsu.edu/hbase/quantum/debrog2.html

OpenStudy (anonymous):

Ok on it

OpenStudy (anonymous):

I think I got it \[\lambda =hc \div pc \] where pc is\[\sqrt{2\times} KE \times m _{0} \times c ^{2}\] and hc is const.